# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v5

@@ Abstract
Text-Based Person Search (TBPS) is a crucial task in the Internet of Things (IoT) domain that enables accurate retrieval of target individuals from large-scale galleries with only given textual caption. For cross-modal TBPS tasks, it is critical to obtain well-distributed representation in the common embedding space to reduce the inter-modal gap. Furthermore, learning detailed image-text correspondences is essential to discriminate similar targets and enable fine-grained search. To address these challenges, we present a simple yet effective method named Sew Calibration and Masked Modeling (SCMM) that calibrates cross-modal representations by learning compact and well-aligned embeddings. SCMM introduces two novel losses for fine-grained cross-modal representations: Sew calibration loss that aligns image and text features based on textual caption quality, and Masked Caption Modeling (MCM) loss that establishes detailed relationships between textual and visual parts. This dual-pronged strategy enhances feature alignment and cross-modal correspondences, enabling accurate distinction of similar individuals while maintaining a streamlined dual-encoder architecture for real-time inference, which is essential for resource-limited sensors and IoT systems. Extensive experiments on three popular TBPS benchmarks demonstrate the superiority of SCMM, achieving 73.81%, 64.25%, and 57.35% Rank-1 accuracy on CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively.
Index Terms:
@@
Text-Based Person Search, Internet of Things (IoT), Person Re-identification, Cross-modal Learning, Metric Learning
I Introduction
With the rapid growth of Internet of Things (IoT) devices and the increasing need for intelligent analysis of visual data, person re-identification (Re-ID) has emerged as a critical task in IoT and multimedia applications [1, 2, 3]. Re-ID aims to identify a target person across non-overlapping camera views, enabling applications such as tracking individuals in crowded scenes and enhancing security measures [4, 5]. However, traditional Re-ID relies on the availability of query images, which may not always be feasible in real-world IoT scenarios. For example, a security guard responding to a witness report describing a suspect’s appearance; in such cases, relying solely on textual descriptions for person retrieval becomes crucial [6, 7]. To address this new scenario, Text-Based Person Search (TBPS) [8] has recently gained increasing interest, which aims to retrieve images of individuals from a large-scale gallery via textual captions as query [9, 10]. In IoT environments, TBPS can be integrated with various sensors and edge server units to enhance real-time processing capabilities and reduce latency. For instance, surveillance cameras  [10] equipped with edge computing units can preprocess video streams and extract textual descriptions of individuals, which are then used for efficient person search across the network. This integration is essential for applications in smart cities and industrial settings  [11], where timely and accurate person identification is crucial for maintaining security and operational efficiency.
@@ I Introduction
Compared to Re-ID, the main task of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can reduce the inter-modal gap, making it easier to locate specific persons with textual captions. Second, detailed correspondence is essential for TBPS. In some cases, fine-grained cross-modal information is necessary to discriminate between two similar persons.
@@
In the TBPS task, numerous approaches have been proposed from two perspectives. Firstly, some methods [12, 13, 14] employ only two encoders and align the two modal representations using a symmetric loss. Although simple, the feature alignment is limited in their methods. The other methods [15, 16] utilize multi-modal models with transformer-based [17] cross-attention to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Secondly, to obtain robust fine-grained features, recent TBPS works have designed multi-level [18, 19, 20, 21], multi-granularity [22, 23] matching strategies, and specific attention modules [10, 24, 25]. These methods rely on the image-text backbone to provide fine-grained features. While these methods provide fine-grained features, they have complex model architectures and costly computations. Furthermore, the fine-grained features they produce are limited and hinder performance boost.
@@
To address these issues, we propose a simple yet effective method called SCMM for calibrating cross-modal representations in TBPS. SCMM consists of only a dual-encoder, making it simple and cost-effective without requiring complex interaction modules or extra multi-level branches, which allows for high-speed inference. In addition, we propose two novel training losses to calibrate cross-modal representations. The first is a Sew calibration loss, which takes the quality of the text description as guidance and aligns features between the textual and visual modalities. It also pushes negative sample pairs apart and pulls positive sample pairs together across the two modalities. Next, we propose a Masking Caption Modeling (MCM) loss to obtain more fine-grained and generic correspondence. This loss uses a masked caption prediction task to establish detailed relationships between text parts and image parts. The operation is implemented through a cross-modal decoder that is discarded at the inference stage, avoiding extra computation costs. To demonstrate the effectiveness of SCMM, we evaluated its performance on three popular benchmarks: CUHK-PEDES [10], ICFG-PEDES [19], and RSTPReID [26]. Our model surpasses the previous state-of-the-art (SOTA) methods and demonstrates impressive performance. Moreover, we conducted extensive experiments to validate each component of our method. Overall, our major contributions can be summarised as follows:
  • We introduce an effective and scalable framework to learn and calibrate cross-modal representations for text-based person search. Our framework utilizes a dual-encoder and an auxiliary cross-modal decoder to achieve efficient and high-speed inference.
  • We propose two novel losses, in which the Sew calibration loss aligns fine-grained features between image and text modalities, as well as the MCM loss establishes detailed relationships between vision modality and textual modality.
  • Extensive experiments demonstrate the superiority of our framework. Our method achieves new state-of-the-art on three popular benchmarks: CUHK-PEDES, ICFG-PEDES, and RSTPReID, which reaches 73.81%, 64.25%, and 57.35% on the Rank1, respectively.
@@
The structure of the remainder of this paper is outlined as follows. Sec. II reviews the related work on text-based person search, metric learning, and masked language modeling. Sec. III introduces the proposed SCMM method, including the Sew Calibration Loss, Masking Caption Modeling Loss, and the overall framework. Sec. IV presents the experimental setup, datasets, implementation details, and comprehensive experimental results, including comparison with state-of-the-art methods, ablation studies on different components, and qualitative analysis with visualization of attention maps and feature distributions. Sec. V concludes the paper with a summary of contributions and future research directions.
II Related Work
II-A Text-Based Person Search
Text-based person search Textis was first introduced by [10], which identifies person images in a gallery only using a textual query. Early works utilized pre-task methods to obtain external cues such as person segmentation [18] and human body landmarks [27]. In recent years, end-to-end frameworks based on attention mechanisms [10, 28, 24, 21, 23, 25] have become prevailing. Cross-modal attention is critical for performing image-text interaction. Existing methods can be broadly classified into attention-explicit and attention-implicit. Specifically, attention-explicit methods [10, 24, 25] design specific attention modules according to multi-granularity and multi-level strategies. For example, NAFS [28] conducts cross-modal alignments over full-scale features with a contextual non-local attention module. CFine [24] utilizes cross-attention for multi-grained global feature learning, it achieves impressive results on three benchmarks with knowledge transfer from the CLIP [29] model. In contrast attention-implicit methods [21, 23] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly. SafaNet [21] introduces a semantic-aligned feature aggregation network. It utilizes a cross-modal parameter-sharing multi-head attention module following the backbone to enhance the extracted image-text representations. However, compared to performing in-depth cross attention with a task-driven approach, the existing methods do not implement enough cross-interaction with generalized performance.
II-B Metric Learning
Initially, metric learning used $L_{2}$ distance as the metric, with the goal is to minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [30] and Triplet Networks [31]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space. L-Softmax [32] introduced the concept of margin on softmax function for the first time. The widely used CosFace [33] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [34] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [35, 36, 11]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS.
II-C Masked Language Modeling
MLM is a highly effective method for pre-training language models [37] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods [38, 15]. For example, the work in [15] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [37] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning.
III Methods
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$. Each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$. Text-based person search aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery with a textual caption. To solve this task, we propose a simple yet effective method, as shown in Fig. 2.
We use ViT [39] and BERT [37] as the image-text encoders in our dual-encoder backbone. The image encoder takes image patches from $I_{i}$ along with a vision classification (CLS) token as input. It outputs an image feature sequence ${v_{i}}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder obtains a text feature sequence ${t_{i}}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following previous works [28, 24].
III-A Sew Calibration Loss with Constraints
In the TBPS task, closely aligning cross-modal representations is crucial for effectively finding specific persons with textual captions. To address the heterogeneity between modalities, we propose a Sew calibration loss that pushes each modality to a common space.

### v7

@@ Abstract
Text-Based Person Search (TBPS) represents a critical engineering application of artificial intelligence (AI) in intelligent surveillance systems where automated person identification using natural language descriptions is essential. Current AI methods face significant limitations in achieving well-aligned cross-modal representations and capturing fine-grained correspondences essential for discriminating similar individuals. The fundamental challenge lies in developing intelligent systems that effectively process heterogeneous data sources where visual and textual features exist in different semantic spaces, creating substantial inter-modal gaps. To address these challenges, we propose Sew Calibration and Masked Modeling (SCMM), a novel AI framework that calibrates cross-modal representations through strategic techniques. SCMM introduces two innovative mechanisms: a sew calibration loss that dynamically aligns image-text features based on textual information quality using adaptive margin learning, and a masked caption modeling loss that establishes detailed cross-modal correspondences through transformer-based masked prediction tasks. The sew calibration component implements quality-aware adaptive constraints, enabling fusion strategies that accommodate varying information richness across modalities. While the masked modeling mechanism facilitates fine-grained feature learning by discovering detailed relationships between visual elements and textual descriptions through cross-modal attention. Our streamlined dual-encoder architecture maintains computational efficiency suitable for real-time applications while achieving superior performance through complementary alignment and correspondence learning strategies. Extensive experiments across three benchmark datasets demonstrate SCMM’s effectiveness, achieving superior Rank1 accuracy on three challenging benchmarks. The framework validates the practical applicability of quality-aware adaptive constraints and fine-grained correspondence modeling for real-world AI engineering applications.
keywords:
@@
Text-Based Person Search, Cross-modal Learning, Person Re-identification, Metric Learning, Engineering Applications, Artificial Intelligence
journal: Engineering Applications of Artificial Intelligence
[…]
[label3]organization=Department of ECE, The University of British Columbia, city=Vancouver, postcode=V6T 1Z4, country=Canada
[…]
[label1]organization=Academy for Engineering &\&& Technology, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[label4]organization=Department of Computer Science, University of Toronto, city=Toronto, postcode=M5S 1A1, country=Canada
[…]
[label2]organization=MEGVII Technology, city=Beijing, postcode=100096, country=China
[…]
[label7]organization=Faculty of Data Science, City University of Macau, city=Macau, postcode=999078, country=China
1 Introduction
Artificial intelligence (AI) applications in engineering have revolutionized modern computer vision and machine learning systems, driven by the urgent need to develop intelligent solutions that can process and integrate heterogeneous data sources for enhanced understanding and comprehensive analysis capabilities [1, 2]. Person re-identification (Re-ID) exemplifies a critical engineering application of AI, requiring sophisticated machine learning algorithms to process visual features across non-overlapping camera views for identifying target individuals in crowded scenes and enhancing security measures in smart city infrastructure [3]. However, traditional Re-ID approaches face significant operational limitations in real-world engineering scenarios when query images are unavailable, necessitating the development of alternative AI-driven solutions [4]. Text-Based Person Search (TBPS) addresses this critical engineering challenge by employing advanced artificial intelligence techniques to process textual descriptions alongside visual galleries, representing a significant advancement in practical AI applications for surveillance and security systems [5]. The implementation of multi-modal processing in TBPS enables the effective integration of rich semantic information from natural language descriptions with visual features from surveillance imagery [6], thereby providing intelligent solutions for Re-ID in scenarios where visual queries are unavailable or insufficient. Consequently, the paradigm shift from traditional visual-only approaches to sophisticated AI-powered cross-modal techniques addresses fundamental limitations in real-world deployment while opening new possibilities for intelligent surveillance and security applications [7].
The effectiveness of AI-powered multi-modal systems depends critically on developing machine learning algorithms that can bridge the semantic gap between heterogeneous modalities while preserving discriminative information essential for accurate person identification [8]. Engineering challenges in TBPS stem from the fundamental differences in feature representations between textual and visual domains, where semantic concepts must be aligned across disparate embedding spaces using advanced AI techniques [9]. State-of-the-art AI architectures have emerged to address these engineering challenges, incorporating transformer-based attention mechanisms and deep learning models for cross-modal interactions to enhance intelligent processing capabilities [10]. Contemporary research in AI engineering applications emphasizes the development of sophisticated alignment strategies that can effectively combine complementary information sources while maintaining computational efficiency suitable for practical deployment in real-world systems [11]. The integration of these advanced AI techniques has proven essential for engineering applications requiring precise person identification with natural language queries, particularly in scenarios where traditional visual-only approaches prove insufficient for accurate target discrimination in intelligent surveillance systems [12].
@@ 1 Introduction
Compared to conventional Re-ID, the main objective of TBPS is to learn fine-grained cross-modal representations between visual and textual modalities [13]. As shown in Fig. 1, cross-modal representations have two key characteristics that contribute to effectively searching for a person from images: closely aligned cross-modal representations and fine-grained information correspondence. First, closely aligned cross-modal representations can effectively reduce the inter-modal gap, making it easier to locate specific persons using textual captions. Second, detailed correspondence is absolutely essential for TBPS, as fine-grained cross-modal information is often necessary to discriminate between similar persons [14].
@@
In the TBPS domain, numerous approaches have been proposed from two distinct perspectives. First, some methods [15, 16] employ only dual encoders and align the two modal representations using symmetric loss functions. Although computationally efficient, these methods suffer from limited feature alignment capabilities. Alternative approaches [17] utilize multi-modal models with transformer-based [18] cross-attention mechanisms to improve cross-modal feature alignment and interaction. However, such methods often require high computational costs due to the exhaustive fusion of all possible image-text pairs during inference, hindering their efficiency in real-time applications. Second, to obtain robust fine-grained features, recent TBPS works have designed multi-level [19, 20, 21, 22], multi-granularity [23, 24] matching strategies, and specialized attention modules [8, 25, 26]. While these methods provide fine-grained features, they suffer from complex model architectures and expensive computations. Furthermore, the fine-grained features they produce are often limited and hinder significant performance improvements.
@@
To address these persistent challenges, we propose a simple yet effective method called SCMM for calibrating cross-modal representations in TBPS. SCMM employs only a dual-encoder architecture, making it computationally efficient without requiring complex interaction modules or additional multi-level branches, which enables high-speed inference suitable for practical applications. Additionally, we propose two novel training losses to effectively calibrate cross-modal representations. The first is a sew calibration loss, which leverages the quality of text descriptions as guidance to align features between textual and visual modalities while pushing negative sample pairs apart and pulling positive sample pairs together across modalities. Furthermore, we propose a Masked Caption Modeling (MCM) loss to obtain more fine-grained and generalizable correspondence. Building upon the principle of masked caption prediction, the loss employs a masked caption prediction task to establish detailed relationships between textual and visual components. The operation is implemented through a cross-modal decoder that is discarded during inference, avoiding additional computational costs. To comprehensively demonstrate SCMM’s effectiveness, we evaluated its performance on three popular benchmarks, where our model surpasses previous state-of-the-art (SOTA) methods and demonstrates impressive performance. Moreover, we conducted extensive experiments to systematically validate each component of our method. Overall, our main contributions can be summarized as follows:
  • 1.
    We propose an efficient and scalable framework for text-based person search, named SCMM, that calibrates cross-modal representations using a dual-encoder architecture and an auxiliary cross-modal decoder. The framework’s design effectively resolves the trade-off between the high computational cost of complex interaction models and the limited alignment capability of simpler methods, achieving superior performance with high inference speed suitable for real-world AI engineering applications.
  • 2.
    We introduce a novel sew calibration loss, which uniquely incorporates a quality-guided adaptive margin mechanism. By dynamically adjusting alignment constraints based on the information density of textual descriptions, our loss function more effectively bridges the inter-modal gap, pushing visual and textual feature distributions into a more compact and well-aligned shared embedding space.
  • 3.
    We propose an innovative MCM loss that establishes detailed cross-modal correspondences through masked prediction tasks and transformer-based cross-modal attention mechanisms. Our method facilitates fine-grained information integration by learning word-level relationships between visual elements and textual descriptions, enabling the framework to distinguish between similar individuals with subtle differences while maintaining computational efficiency through training-only decoder utilization.
  • 4.
    We conduct extensive experiments on three challenging benchmarks: CUHK-PEDES, ICFG-PEDES, and RSTPReID. SCMM demonstrates superiority by achieving new state-of-the-art Rank1 accuracies of $73.81\%$, $64.25\%$, and $57.35\%$ respectively. The establishment of new performance benchmarks not only validates the effectiveness of our fusion strategies for text-based person search, but also provides valuable baselines for future research in cross-modal AI applications.
The structure of the remainder of this paper is outlined as follows. Sec. 2 reviews the related work on text-based person search, metric learning, and masked language modeling. Sec. 3 introduces the proposed SCMM method, including the sew calibration loss, masked caption modeling loss, and the overall framework. Sec. 4 presents the experimental setup, datasets, implementation details, and comprehensive experimental results, including comparison with state-of-the-art methods, ablation studies on different components, and qualitative analysis with visualization of attention maps and feature distributions. Sec. 5 concludes the paper with a summary of contributions and future research.
2 Related Work
2.1 Text-Based Person Search
Text-based person search was first introduced by [8], which identifies person images in a gallery using only textual queries. Fundamentally, the task represents a critical application of multi-modal fusion, where the integration of heterogeneous data sources—visual and textual modalities—becomes essential for enhanced understanding and analysis capabilities. Early works utilized pre-processing methods to obtain external cues such as person segmentation [19] and human body landmarks [27]. The evolution towards sophisticated fusion architectures has become crucial, as multi-view fusion combining complementary information from multiple sources shows significant promise for enhancing prediction accuracy and robustness in cross-modal tasks [10].
In recent years, end-to-end frameworks based on attention mechanisms [22, 24, 26] have become prevalent, addressing the fundamental challenge of semantic mismatch between features of different modalities. Cross-modal attention is critical for performing image-text interaction, where advanced fusion strategies must consider the unique attributes of cross-modal matching tasks [3]. Existing methods can be broadly classified into attention-explicit and attention-implicit approaches. Attention-explicit methods [2, 26] design specific attention modules according to multi-granularity and multi-level strategies, implementing early fusion techniques that combine features at the input level. For example, NAFS [14] conducts cross-modal alignments over full-scale features with a contextual non-local attention module, while CFine [25] utilizes cross-attention for multi-grained global feature learning, achieving impressive results through knowledge transfer from the CLIP [28] model.
In contrast, attention-implicit methods [22, 24] utilize transformer-based models with shared parameters to align cross-modal semantics implicitly, representing late fusion approaches that combine decisions from individual modality-specific models [10]. However, compared to performing in-depth cross-attention with task-driven approaches, existing methods face challenges in addressing information asymmetry problems and do not implement sufficient cross-interaction with generalized performance, highlighting the need for more sophisticated fusion techniques that enhance feature integration and establish deeper connections between global structure and local details [3].
2.2 Metric Learning
Initially, metric learning used $L_{2}$ distance as the metric, with the goal is to minimize the $L_{2}$ distance between samples of the same class. Some $L_{2}$-based metric learning methods include Siamese Networks [29] and Triplet Networks [30]. With the development of deep learning, researchers started using softmax-based loss functions in metric learning to learn a more discriminative distance metric. Additionally, increasing the margin between classes is an intuitive approach to learn a better metric space [3]. L-Softmax [31] introduced the concept of margin on softmax function for the first time. The widely used CosFace [32] proposed large-margin cosine loss to learn highly discriminative deep features for face recognition. Circle loss [33] proposed a simple loss based on a unified loss for metric learning and classification. Recently, some works introduced adaptive margin into marginal loss [2]. They usually learn image quality implicitly and adjust the margin accordingly. In single-modal representation learning, they usually give large margins to high-quality samples for hard mining. Compared to them, we try to solve a cross-modal matching problem where samples from two modalities have different information volume. We give greater tolerance to less informative samples in TBPS.
However, existing metric learning approaches face significant limitations when applied to cross-modal scenarios. First, traditional methods primarily focus on single-modal optimization and fail to adequately address the heterogeneity gap between different modalities [34]. Second, fixed margin constraints cannot adaptively handle the varying information density across modalities, leading to suboptimal alignment performance. Finally, most current approaches lack fine-grained correspondence modeling capabilities essential for distinguishing between similar cross-modal samples [4].
2.3 Masked Language Modeling
MLM is a highly effective method for pre-training language models [35] by randomly selecting a certain percentage of words from the input sentence and then predicting the masked words based on the context of other words. Many cross-modal pre-training models have utilized MLM in their methods. For example, the work in [17] combines MLM with contrastive loss in the framework, which achieved impressive performance in their cross-modal tasks. The success of MLM in BERT [36] has proven its ability to adapt well to various downstream tasks, leading to generalized performance. In our fine-grained framework, the task-driven decoder utilizing masked caption modeling can facilitate generic cross-modal learning.
However, existing MLM-based cross-modal approaches face several limitations when applied to fine-grained person search tasks. First, traditional MLM methods primarily focus on intra-modal language understanding and lack explicit cross-modal interaction mechanisms, limiting their ability to establish detailed visual-textual correspondences [6]. Second, most existing approaches apply MLM uniformly without considering the varying information density and quality of textual descriptions, leading to suboptimal learning efficiency. Finally, current methods often require complex multi-stage training procedures or heavy computational overhead during inference, hindering their practical deployment in real-time applications [23].
3 Proposed Methods
Formally, given a set of images with corresponding captions, denoted as $X=\{(I_{i},T_{i})\}^{N}_{i=1}$, where each image $I_{i}$ and its description text $T_{i}$ is associated with a person ID $y_{i}$. The text-based person search task aims to retrieve the most relevant Rank $k$ (e.g., $k=1,5,10$) person images efficiently from a large-scale gallery using textual queries. As an engineering application of AI, it requires developing machine learning algorithms that can bridge the semantic gap between visual and textual modalities while maintaining discriminative capabilities for accurate person identification in real-world scenarios. The fundamental engineering challenge lies in designing AI systems that can align heterogeneous feature representations across modalities, where visual features capture spatial and appearance characteristics while textual descriptions encode semantic attributes and contextual information. Effective strategies must address both global alignment between modalities and fine-grained correspondences that enable discrimination between similar individuals in practical applications.
@@ 3 Proposed Methods
To address these engineering challenges in AI applications, we propose SCMM, a streamlined yet effective machine learning framework as illustrated in Fig. 2. Our approach implements a dual-encoder architecture using state-of-the-art AI models: ViT [37] and BERT [36] as backbone networks for visual and textual feature extraction respectively. The image encoder processes image patches from $I_{i}$ along with a vision classification (CLS) token, outputting an image feature sequence $\{v_{i}\}$ and a vision CLS token embedding $v_{i}^{c}$. Similarly, the text encoder generates a text feature sequence $\{t_{i}\}$ and a text CLS token embedding $t_{i}^{c}$ for caption $T_{i}$, following established AI engineering practices [14, 25]. Our framework integrates two complementary fusion mechanisms to calibrate cross-modal representations: the sew calibration loss with adaptive margin constraints for global alignment, and the masked caption modeling loss for establishing detailed cross-modal correspondences. The following subsections elaborate on each component. We first introduce the sew calibration loss that dynamically aligns cross-modal features based on textual information quality using machine learning techniques. Subsequently, we present the masked caption modeling mechanism that facilitates fine-grained information integration through transformer-based cross-modal attention and masked prediction tasks. The total loss function combines these complementary AI objectives to jointly optimize global cross-modal alignment and fine-grained correspondence learning, enabling the framework to achieve superior performance while maintaining computational efficiency suitable for real-world engineering applications through the streamlined dual-encoder architecture.
3.1 Sew Calibration Loss with Constraints